# Structural Properties of Entropic Vectors and Stability of the Ingleton Inequality

### v1

@@ 1 Introduction
At a high level, the proofs of Theorems A and B follow a similar pattern. We first apply Corollary E, and then use standard Shannon–type inequalities on the original variables together with the auxiliary variable $W$; see the full argument in Section 4. The proofs presented in this paper are entirely conventional in the sense that we explicitly write down the necessary entropy inequalities. However, the same arguments can be generated or mechanically verified with computer assistance. In practice, one may apply Corollary E in a suitable form, take the properties (a)–(c) for the resulting auxiliary variable $W$, include all basic Shannon-type inequalities involving the variables under consideration (including $W$), and then use a linear-programming solver to deduce the desired conclusion. Since the bounds in Corollary E hold only up to $O(\sqrt{\varepsilon})$, the resulting inequality must likewise be understood up to an error term of order $O(\sqrt{\varepsilon})$.
Furthermore, we combine Lemma D with symmetry considerations and obtain two additional results. One of these results demonstrates somewhat subtler properties of tuples of random variables. If the six-tuple $(X,Y,A,B,A^{\prime},B^{\prime})$ of random variables satisfies the bounds
$\max\{I(A:X|Y),I(A:Y|X),I(B:X|Y),I(B:Y|X)\}\leq\varepsilon\cdot I(X:Y),$
then the following alternative holds: either a stronger version of the Ingleton inequality for $(X,Y,A,B)$ holds
$I(X:Y)\textrm{ is much less than }I(X:Y|A)+I(X:Y|B)+I(A:B)$
or a slightly weaker version of the Ingleton inequality for $(X,Y,A^{\prime},B^{\prime})$ holds
$I(X:Y)\ \textrm{ is not much bigger than }I(X:Y|A^{\prime})+I(X:Y|B^{\prime})+I(A^{\prime}:B^{\prime}).$
Here is the exact statement.
Theorem F (name=,restate=thsubtler).
If the joint distribution $(X,Y,A,B,A^{\prime},B^{\prime})$ satisfies
$\max\{I(A:X|Y),I(A:Y|X),I(B:X|Y),I(B:Y|X)\}\leq\varepsilon\cdot I(X:Y)$
Then
  1. (a)
    The following alternative holds: for every threshold $0<t<1$ either we have a stronger version of Ingleton’s inequality for $(A,B,X,Y)$
    or a weaker version of Ingleton’s inequality for $(A^{\prime},B^{\prime},X,Y)$
    $I(X:Y)\leq I(X:Y|A^{\prime})+I(X:Y|B^{\prime})+I(A^{\prime}:B^{\prime})+\big(O(\sqrt{\varepsilon})+t\big)I(X:Y)$
  2. (b)
    Besides, we have the following inequality for six random variables
    $\;\;+I(X:Y|A^{\prime})+I(X:Y|B^{\prime})+I(A^{\prime}:B^{\prime})$
Note that the assumptions in the theorem above do not involve variables $A^{\prime},B^{\prime}$ in any capacity, so that the conclusion of the theorem remains true for arbitrary $A^{\prime}$, $B^{\prime}$.
Part (b) of Theorem F probably cannot be proven with more direct approaches (without applying explicitly or implicitly infinitely many instances of the Copy Lemma). This theorem does not seem to follow from any previously known results. We prove it in Section 5.
Another result in which we combine Lemma D with symmetry considerations is Theorem C. In this theorem we assume that four conditional mutual information terms are of order $\varepsilon$. Under this assumption, we obtain a version of the Ingleton bound with an error term of order $O(\varepsilon\log\varepsilon^{-1})$ (instead of $O(\sqrt{\varepsilon})$ in the previous theorems). The proof of Theorem C is presented in Section 6.
@@ 2.1 Notation
In what follows we use the following notation and conventions. We denote $n]:=\left\{0,1,\dots,n-1\right\}$ and $2^{[n]}$ will stand for the power set of $\left\{0,1,\dots,n-1\right\}$. For a collection of jointly distributed random variables $(X_{0},\dots,X_{n-1})$ and $J\in 2^{[n]}$, we denote by $X_{J}$ the joint random variable corresponding to the subcollection $(X_{i}\;{\bm{:}}\;i\in J)$. The entropy of a random variable $X$ (respectively, the conditional entropy of $X$ given $Y$) will be denoted $H(X)$ (and, respectively, $H(X|Y)$). We use the standard notation for the mutual information in pairs and triples of random variables,
@@ 2.2 Tropicalization/Tensorization
$\operatorname{\mathsf{dist}}(\mathbf{X}^{(k)}\otimes\mathbf{X}^{(l)},\mathbf{X}^{(k+l)})\leq C\cdot\varphi(k+l)$
where $\operatorname{\mathsf{dist}}$ stands for the entropic distance between $n$-tuples of random variables and $\varphi$ is an admissible (sufficiently slowly growing) function, and $C$ is some positive constant that may depend on $\mathbf{X}$. For example, we can take $\varphi(t):=t^{3/4}$; see [MP20] for details.
@@
From now on, when we say a (tuple of) random variables we mean tropical (tuple of) random variables. Entropy of such a tropical random variable is defined as the limit of scaled entropies of individual members of the sequence.
Two observations are in order. First of all, any linear inequality that is valid for entropies for all tropical random variables implies a similar inequality for entropies of conventional random variables. Secondly, if every component of a tropical tuple obeys a given linear inequality with a sublinear error term, e.g., up to $O(1)$, then the tropical tuple itself satisfies this inequality exactly (with no error). Thus, these “tropical” objects can be interpreted as a helpful tool to study properties of the more conventional objects such as probability distributions.
@@ 2.3 Symmetries in distribution and tropical random variables
Thus, we symmetrize the distribution, while changing any $(A,B)$-symmetric entropic expressions by $O(1)$, uniformly with respect to the data. Given a tropical tuple we can perform the above procedure elementwise in the sequence of tuples, thereby obtaining exact equalities for the symmetric entropic expressions. ∎
Lemma 3 implies immediately the following corollary.
Corollary 4.
Let $(X_{0},\dots,X_{n-1},A,B)$ be a tropical tuple such that for any $J\in 2^{[n]}$ one has $H(X_{J},A)=H(X_{J},B).$ Then there exists another tropical tuple $(X_{0},\dots,X_{n-1},\bar{A},\bar{B})$ such that
  • the probability distribution of the new tuple is symmetric with respect to the transposition of $\bar{A}$ and $\bar{B}$, and
  • for any $J\in 2^{[n]}$ one has
    $H(X_{J},\bar{A})=H(X_{J},A),$
    $H(X_{J},\bar{B})=H(X_{J},B),$
    $H(X_{J},\bar{A},\bar{B})=H(X_{J},A,B).$
Thus, if the entropy profile of a tuple is symmetric, we can replace it with a tuple with the same entropy profile and a symmetric distribution.
Remark 5.
Let ${\cal L}_{i}(X,Y,A,B)$, $i=1,2$, be linear combinations of entropies involving $X,Y,A,B$. Assume also that these ${\cal L}_{i}$ are symmetric (as formal expressions) with respect to the transposition of the symbols $A$ and $B$, e.g. the expression from the Ingleton inequality:
$\begin{array}[]{l}I(A:B|X)+I(A:B|Y)+I(X:Y)-I(A:B)\\ =H(A,X)+H(B,X)-H(A,B,X)+H(A,Y)+H(B,Y)-H(A,B,Y)\\ \phantom{=}-H(X,Y)-H(A)-H(B)+H(A,B)\end{array}$
In this setting, if we can prove a statement of the form
${\cal L}_{1}(X,Y,A,B)=0\Rightarrow{\cal L}_{2}(X,Y,A,B)\geq 0$
for the distributions symmetric with respect to the transposition of $A$ and $B$, then this implication is true for all distributions. Indeed, given an arbitrary distribution $(X,Y,A,B)$, we can apply Lemma 3 and construct its symmetrized version. If ${\cal L}_{1}(X,Y,A,B)=0$ for the original distribution, the same equality holds for the symmetrized one (here we use symmetry of ${\cal L}_{1}$). So we can conclude that ${\cal L}_{2}(X,Y,A,B)\geq 0$ for the symmetrized distribution. Finally, we use symmetry of ${\cal L}_{2}$ and conclude that the same inequality holds for the original distribution.
This observation allows us in some cases to assume without loss of generality that the random variables under consideration have the property of symmetry.
@@ 2.4 Copy lemma
$(X_{i},Y_{j},A^{\prime}\;{\bm{:}}\;i=0,\ldots,k-1;\;j=0,\dots,\ell-1)$
@@
The entropy profile of the extended tuple $(X_{[k]},Y_{[\ell]},A,A^{\prime})$ is not totally defined by the entropies of the original distribution $(X_{[k]},Y_{[\ell]},A)$. However, some components of the new entropy profile are determined by the definition of $A^{\prime}$: for all $I\subset[k]$ and $J\subset[\ell]$
@@ 2.5 Ahlswede–Körner reduction
The Ahlswede–Körner reduction, [AK06, AK77], modifies a given tropical $n$-tuple so that some entropies in the new tuple vanish, while some (combinations of) other entropies stay the same.
@@
Let us consider a tropical $n$-tuple $(X_{i}:i=0,\dots,n-1)$ and a subcollection $(X_{i}:i=0,\dots,k-1)$ of it, where $k\leq n$. Then there exists an $(n+1)$-tuple $(X_{0},\dots,X_{n-1},X^{\prime}_{n-1})$, such that
@@ 3 Main construction
In this section we prove our main technical results, Lemma D and Corollary E, which are repeated below verbatim. They are systematically applied in all proofs of the article. Assume we are given a jointly distributed triple $(X,Y,A)$ or, possibly, a longer joint distribution $(X,Y,A,B,C_{0},\ldots,C_{n-1})$ extending $(X,Y,A)$. For each integer $k\geq 0$, we extend this tuple with a new variable $W_{k}$ satisfying the properties stated in the lemma.
@@ 3.1 The main construction
First of all, we apply the Copy Lemma $k$ times to $A$ under $(X,Y)$ and obtain new variables $A^{\prime}_{1},\ldots,A^{\prime}_{k}$ such that
  • for each $i=1,\ldots,k$ the conditional distribution of $A^{\prime}_{i}$ given $(X,Y)$ is the same as the conditional distribution of $A$;
@@
Further, we apply the Ahlswede–Körner reduction to $\tilde{A}_{k}$ under $(A,B,C_{0},\ldots,C_{n-1})$ and denote the result by
@@
3.2 Proof of Lemma D
We use $W_{[k]}$ constructed in the previous section to prove the lemma. To prove assertion (a), we note that by the Copy Lemma, the variable $\tilde{A}_{k}$ is independent of $(A,B,C_{0},\ldots,C_{n-1})$ conditioned on $(X,Y)$. The same holds for $W_{k}=\operatorname{\mathsf{AK}}(\tilde{A}_{k}|X,Y,A,B,C_{0},\dots,C_{n-1})$, by the definition of the Ahlswede–Körner reduction. Thus we have
$H(W_{k}|X,Y)=H(W_{k}|X,Y,A,B,C_{0},\dots,C_{n-1})=0$
To prove assertion (b) it suffices to observe that the conditional distributions of $(X,Y)$ given $(A^{\prime}_{1},\ldots,A^{\prime}_{k+1})$ and given $(A^{\prime}_{1},\ldots,A^{\prime}_{k},A)$ are identical.
To prove (c), we observe that
$H(W_{k}|X)$ $=I(\tilde{A}_{k}:X,Y,A,B,C_{0},\ldots,C_{n-1}|X)$

### v2

@@ 1.4 Methods and ideas
At a high level, the proofs of Theorems A and B follow a similar pattern. We first apply Corollary F, and then use standard Shannon–type inequalities on the original variables together with the auxiliary variable $W$; see the full argument in Section 4. The proofs presented in this paper are entirely conventional in the sense that we explicitly write down the necessary entropy inequalities. However, the same arguments can be generated or mechanically verified with computer assistance. In practice, one may apply Corollary F in a suitable form, take the properties (a)–(c) for the resulting auxiliary variable $W$, include all basic Shannon-type inequalities involving the variables under consideration (including $W$), and then use a linear-programming solver to deduce the desired conclusion. Since the bounds in Corollary F hold only up to $O(\sqrt{\varepsilon})$, the resulting inequality must likewise be understood up to an error term of order $O(\sqrt{\varepsilon})$.
In the proof of Theorem C (see Section 6) and Theorem D (see Section 5) we combine Lemma E with symmetry considerations.
@@ 2.1 Notation
In what follows we use the following notation and conventions. We denote $n]:=\left\{0,1,\dots,n-1\right\}$ and $2^{[n]}$ will stand for the power set of $\left\{0,1,\dots,n-1\right\}$. In this paper, when we talk about random variables, we always mean finitely supported random variable. For a collection of jointly distributed random variables $(X_{0},\dots,X_{n-1})$ and $J\in 2^{[n]}$, we denote by $X_{J}$ the joint random variable corresponding to the subcollection $(X_{i}\;{\bm{:}}\;i\in J)$. The entropy of a random variable $X$ will be denoted $H(X)$. We use the standard notation for the conditioned entropy, the mutual information in pairs and triples of random variables,
$H(X|Y):=H(XY)-H(Y)$
@@ 2.2 Tropicalization/Tensorization
$\operatorname{\mathsf{dist}}(\mathbf{X}^{(k)}\otimes\mathbf{X}^{(l)},\mathbf{X}^{(k+l)})\leq C\cdot\psi(k+l)$
where $\operatorname{\mathsf{dist}}$ stands for the entropic distance between $n$-tuples of random variables and $\psi$ is an admissible [^2] A function $\psi$ is admissible if it is monotone and sufficiently slowly growing, more specifically $\int_{1}^{\infty}(\psi(t)/t^{2})\mathrm{d}t<\infty$ function, and $C$ is some positive constant that may depend on the sequence $(\mathbf{X}^{(k)})_{k\in\mathbb{N}}$. For example, we can take $\psi(t):=t^{3/4}$ and fix this choice in this article; see [MP20] for details.
@@
From now on, when we say a (tuple of) random variables we mean tropical (tuple of) random variables. Entropy of such a tropical random variable is defined as the limit of scaled entropies of individual members of the sequence. Other properties of tropical tuples, such as symmetry, are to be understood in the sense that they apply to the tropical sequence element-wise for sufficiently large $n$
Remark 5.
Two observations are in order.
  1. (i)
    First of all, any linear inequality that is valid for entropies for all tropical random variables implies the same inequality for entropies of conventional random variables.
  2. (ii)
    Secondly, if every component of a tropical tuple obeys a given linear inequality with a sublinear error term, e.g., up to a constant independent of $k$, then the tropical tuple itself satisfies this inequality exactly (with no error).
Thus, these “tropical” objects can be interpreted as a helpful tool to study properties of the more conventional objects such as probability distributions.
@@ 2.3 Symmetries in distribution and tropical random variables
Thus, we symmetrize the distribution, while changing any $(A,B)$-symmetric entropic expressions by $O(1)$, uniformly with respect to the data. Given a tropical tuple we can perform the above procedure elementwise in the sequence of tuples, thereby obtaining exact equalities for the symmetric entropic expressions, see Remark 5. ∎
@@ 2.4 Copy lemma
$(X_{i},Y_{j},A,A^{\prime}\;{\bm{:}}\;i=0,\ldots,k-1;\;j=0,\dots,\ell-1)$
@@
The entropy profile of the extended tuple $(X_{[k]},Y_{[\ell]},A,A^{\prime})$ is not totally determined by the entropies of the original distribution $(X_{[k]},Y_{[\ell]},A)$. However, some components of the new entropy profile are determined by the definition of $A^{\prime}$: for all $I\subset[k]$ and $J\subset[\ell]$
@@ 2.5 Ahlswede–Körner reduction
The Ahlswede–Körner reduction modifies a given tropical $n$-tuple so that some entropies in the new tuple vanish, while some (combinations of) other entropies stay the same.
@@
Let us consider a tropical $n$-tuple $(X_{i}:i=0,\dots,n-1)$ and a subcollection $(X_{i}:i=0,\dots,k-1)$ of it, where $k<n$. Then there exists an $(n+1)$-tuple $(X_{0},\dots,X_{n-1},X^{\prime}_{n-1})$, such that
@@
Remark 9.
The claim $H(X_{J}|X^{\prime}_{n-1})=H(X_{J}|X_{n-1})$ in the Ahlswede–Körner reduction can be reformulated as $I(X_{J}:X^{\prime}_{n-1})=I(X_{J}:X_{n-1})$.
Remark 10.
The extension of the distribution guaranteed by the Ahlswede–Körner reduction is not unique, i.e., there may exist different (nonisomorphic) extensions of $(X_{i}:i=0,\dots,n-1)$ satisfying the conditions of Lemma 8. We abuse notation and write $X^{\prime}_{n-1}:=\operatorname{\mathsf{AK}}(X_{n-1}\mid X_{0},\dots,X_{k-1})$, meaning that any such extension may be chosen.
Essentially Lemma 8 claims that the variable $X_{n-1}$ can be “reduced” to a new random variable $X^{\prime}_{n-1}$, that will be functionally dependent on $X_{[k]}$, while all the quantities $I(X_{J}:X_{n-1})$, $J\in 2^{[k]}$, will remain preserved after the reduction. In Fig. 1 we illustrate this reduction for the case $k=n=2$. This statement is not literally true for untropicalized tuple, but only hold up to an additive error, which is sublinear with respect to $H(X_{[n]})$. Recall that tropical tuple is a growing sequence of the classical tuples of random variables such that entropic quantities of individual members of the sequence grow quasi-linearly. Entropic quantity evaluated on the tropical tuple is the limit of the corresponding quantities of individual tuples normalized by the index in the sequence. Thus after tropicalization sublinear errors vanish.
[…]
Figure 1: Diagrams representing the Ahlswede–Körner reduction for $k=2$. The reduced variable $X_{2}^{\prime}=\operatorname{\mathsf{AK}}(X_{n-1}\mid X_{0},X_{1})$ has the same mutual information with $X_{0}$, $X_{1}$, and $(X_{0},X_{1})$ as the original variable $X_{2}$; at the same time, $H(X_{2}^{\prime}|X_{0},X_{1})$ vanishes.
History of the Ahlswede–Körner reduction.
This technical tool goes back to the work of Ahlswede and Körner in the 1970s [AK06, AK77]. For readers interested in a proof of this lemma, we can point to several sources. For the case $k=2$ and $n=3$, the statement is proven in [MMRV02, Lemma 5], and the argument extends straightforwardly to arbitrary $k$ and $n$. A similar argument can be found in [MP19, Theorem 3.1]. The latter article, however, uses somewhat different terminology.
The lemma can also be obtained as a special case of more general techniques. Namely, one may first introduce a copy $X^{\prime}_{n-1}$ of $X_{n-1}$ (preserving its joint distribution with $(X_{0},\dots,X_{k-1})$), and then perform tightening as in [MC16, Lemma 3] (with $L=\{X^{\prime}_{n-1}\}$ and $t=H(X_{n-1}\mid X_{0},\dots,X_{k-1})$). The technically most involved part of this argument is established in [Mat07a, Theorem 3], which can be viewed as a generalization of the Slepian–Wolf theorem [SW73]. An overview of the tightening technique and its applications can be found in [Csi25].
@@ 3 Main construction
In this section we prove our main technical results, Lemma E and Corollary F, which are repeated below verbatim. They are systematically applied in all proofs of the article. Assume we are given a jointly distributed triple $(X,Y,A)$ or, possibly, a longer joint distribution $(X,Y,A,B,C_{0},\ldots,C_{n-1})$ extending $(X,Y,A)$. For each integer $k\geq 0$, we extend this tuple with a new variable $W_{k}$ satisfying the properties stated in the lemma.
@@ 3.1 The main construction
First of all, we apply the Copy Lemma $k$ times to $A$ under $(X,Y)$ and obtain new variables $A^{\prime}_{1},\ldots,A^{\prime}_{k}$ such that for each $i=1,\ldots,k$
  • the new variable $A^{\prime}_{i}$ has the same distribution as $A$ conditional on $(X,Y)$;
  • conditional on $(X,Y)$, the new variable $A^{\prime}_{i}$ is independent of $A^{\prime}_{1},\ldots,A^{\prime}_{i-1},A,B,C_{0},\dots,C_{n-1}$.
Observe that
$I(A^{\prime}_{1},\ldots,A^{\prime}_{k}$ $:A,B,C_{0},\dots,C_{n-1}|X,Y)$
$=\sum\limits_{i=1}^{k}I(A^{\prime}_{i}:A,B,C_{0},\dots,C_{n-1}|X,Y,A^{\prime}_{1},\ldots,A^{\prime}_{i-1})$
$\leq\sum\limits_{i=1}^{k}I(A^{\prime}_{i}:A^{\prime}_{1},\ldots,A^{\prime}_{i-1},A,B,C_{0},\dots,C_{n-1}|X,Y)=0$
(the first step is the chain rule for the mutual information; the second step is an instance of the Shannon-type inequality $I(U:V|W)\leq I(U:V,W)$; the last equality follows from the definition of $A^{\prime}_{i}$ in the Copy Lemma). Thus, we conclude that
@@
Further, we apply the Ahlswede–Körner reduction to $\tilde{A}_{k}$ under the tuple $(A,B,C_{0},\ldots,C_{n-1})$ and denote the result by
@@
3.2 Proof of Lemma E
We recall that we define each variable $A^{\prime}_{i+1}$ with help of the Copy Lemma so that
  1. (a)
    it has the same distribution as $A$ conditional on $(X,Y)$;
  2. (b)
    $A^{\prime}_{i+1}$ and $(A^{\prime}_{1},\ldots,A^{\prime}_{i},A,B,C_{0},\ldots,C_{n-1})$ are independent conditional on $(X,Y)$.
Claim 11.
For each $k$
$(A,B,C_{0},\ldots,C_{n-1})\to(X,Y)\to(A^{\prime}_{1},\ldots,A^{\prime}_{k})$
is a Markov chain.
This lemma follows easily from the definition of the quantities $A^{\prime}_{i}$. For completeness of exposition, we give a detailed derivation in terms of information inequalities.
Proof of the claim:.
The argument is based on the property (b) above with $i=1,\ldots,k-1$. We use the chain rule, and then apply $i$ times inequalities having the form (23) and obtain
$I(A,B,C_{0},\ldots,C_{n-1}:A^{\prime}_{1},\ldots,A^{\prime}_{i}|X,Y)$ $=\sum\limits_{j=0}^{i}I(A,B,C_{0},\ldots,C_{n-1}:A^{\prime}_{j}|X,Y,A^{\prime}_{1},\ldots,A^{\prime}_{j-1})$
$\leq\sum\limits_{j=0}^{i}I(A^{\prime}_{1},\ldots,A^{\prime}_{j-1},A,B,C_{0},\ldots,C_{n-1}:A^{\prime}_{j}|X,Y)$
$=0,$
and the claim is proven. ∎
Now we can prove assertion (a) of Lemma E. Claim 11 implies
$I(\tilde{A}_{k}:A,B,C_{0},\ldots,C_{n-1},X,Y)$ $\leq I(\tilde{A}_{k}:X,Y)+I(\tilde{A}_{k}:A,B,C_{0},\ldots,C_{n-1}|X,Y)$
$=I(\tilde{A}_{k}:X,Y).$
From the definition of the Ahlswede–Körner reduction (see Remark 9) it follows
$I(W_{k}:X,Y,A,B,C_{0},\ldots,C_{n-1})$ $=I(\tilde{A}_{k}:X,Y,A,B,C_{0},\ldots,C_{n-1})$
$\leq I(\tilde{A}_{k}:X,Y)=I(W_{k}:X,Y).$
Therefore, with the chain rule we obtain
$I(W_{k}:A,B,C_{0},\ldots,C_{n-1}|X,Y)$ $=I(W_{k}:X,Y,A,B,C_{0},\ldots,C_{n-1})-I(W_{k}:X,Y)$
$=0.$
Hence,
$H(W_{k}|X,Y)$ $=H(W_{k}|X,Y,A,B,C_{0},\ldots,C_{n-1})+I(W_{k}:A,B,C_{0},\ldots,C_{n-1}|X,Y)$
$=H(W_{k}|X,Y,A,B,C_{0},\ldots,C_{n-1})=0$
(the last equality follows from the Ahlswede–Körner reduction).
To prove assertion (b) of Lemma E, we observe that the following conditional distributions are all identical:
  1. (i)
    distribution of $A$ given the value of $(X,Y,A^{\prime}_{1},\ldots,A^{\prime}_{i})$
  2. (ii)
    distribution of $A$ given the value of $(X,Y)$
    (this distribution is equivalent to the previous one due to the Markov property)
  3. (iii)
    distribution of $A^{\prime}_{i+1}$ given the value of $(X,Y)$
    (equivalent to the previous one due to the property (a) above)
  4. (iv)
    distribution of $A^{\prime}_{i+1}$ given the value of $(X,Y,A^{\prime}_{1},\ldots,A^{\prime}_{i})$
    (equivalent to the previous one due to the property (b) above).
Comparing (i) and (iv), we conclude that the conditional distributions
[…]
and
[…]
are the same. It follows that
$\left\{\begin{array}[]{l}H(X,Y|\tilde{A}_{i+1})=H(X,Y|\tilde{A}_{i},A)=H(X,Y,A|\tilde{A}_{i})-H(A|\tilde{A}_{i}),\\ H(X|\tilde{A}_{i+1})=H(X|\tilde{A}_{i},A)=H(X,A|\tilde{A}_{i})-H(A|\tilde{A}_{i}),\\ H(Y|\tilde{A}_{i+1})=H(Y|\tilde{A}_{i},A)=H(Y,A|\tilde{A}_{i})-H(A|\tilde{A}_{i}).\end{array}\right$.
The Ahlswede–Körner reduction guarantees that
[…]
and
$\left\{\begin{array}[]{l}H(X,Y|\tilde{A}_{i})=H(X,Y|W_{i}),\\ H(X|\tilde{A}_{i})=H(X|W_{i}),\\ H(Y|\tilde{A}_{i})=H(Y|W_{i}),\\ H(X,Y,A|\tilde{A}_{i})=H(X,Y,A|W_{i}),\\ H(X,A|\tilde{A}_{i})=H(X,A|W_{i}),\\ H(Y,A|\tilde{A}_{i})=H(Y,A|W_{i}),\\ H(A|\tilde{A}_{i})=H(A|W_{i}).\end{array}\right$.
We combine these observations and obtain
$I(X:Y|W_{i+1})$ $=H(X|W_{i+1})+H(Y|W_{i+1})-H(X,Y|W_{i+1})$
$=H(X|\tilde{A}_{i+1})+H(Y|\tilde{A}_{i+1})-H(X,Y|\tilde{A}_{i+1})$
$=H(X|\tilde{A}_{i},A)+H(Y|\tilde{A}_{i},A)-H(X,Y|\tilde{A}_{i},A)$
$=H(X,A|\tilde{A}_{i})+H(Y,A|\tilde{A}_{i})-H(X,Y,A|\tilde{A}_{i})-H(A|\tilde{A}_{i})$
$=H(X,A|W_{i})+H(Y,A|W_{i})-H(X,Y,A|W_{i})-H(A|W_{i})$
$=I(X:Y|W_{i},A),$
and (b) is proven.
To prove (c) of Lemma E, we observe that
$H(W_{k}|X)$ $=H(W_{k},X)-H(X)$
$=H(X|W_{k})+H(W_{k})-H(X)$
$=H(X|\tilde{A}_{k})+I(\tilde{A}_{k}:X,Y,A,B,C_{0},\ldots,C_{n-1})-H(X)$